The following fractions represent just three different numbers. Separate them into three groups of equivalent fractions, by changing each one to its simplest form.
3/15

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3 / 15 = (1× 3) / (5 × 3)
= 1/5

Thanks